# Discriminating image representations with principal distortions

### v1

@@ 3.2 Principal distortions for comparing multiple image representations
$\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}=\operatorname*{arg\,max}_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}\sum_{n=1}^{N}\sum_{m=1}^{N}m_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}^{2}({\bm{I}}_{A_{n}},{\bm{I}}_{A_{m}}).$ (4)
@@
$\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}=\operatorname*{arg\,max}_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}\sum_{n=1}^{N}\left|\log\frac{d_{A_{n}}({\bm{\epsilon}})}{d_{A_{n}}({\bm{\epsilon}}^{\prime})}-\frac{1}{N}\sum_{m=1}^{N}\log\frac{d_{A_{m}}({\bm{\epsilon}})}{d_{A_{m}}({\bm{\epsilon}}^{\prime})}\right|^{2}$
@@
Note that there are several other natural extensions of generalized-eigendistortions when considering $N>2$ models. For example, for any $p\geq 1$, one can choose distortions ${\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}$ that maximize the sum of the $p$th power of all pairwise distances, which amounts to replacing $m_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}^{2}({\bm{I}}_{A_{n}},{\bm{I}}_{A_{m}})$ in Equation 4 with $m_{{\bm{\epsilon}},{\bm{\epsilon}}^{\prime}}^{p}({\bm{I}}_{A_{n}},{\bm{I}}_{A_{m}})$. Here we focus on the case of maximizing the variance ($p=2$) and leave an exploration of higher-order moments to future work
@@ 4 Experimental results
As a demonstration of our method, we generated principal distortions for computational models previously proposed to capture aspects of the human visual system. All models were implemented in PyTorch (Ansel et al., 2024) and simulations were performed on NVIDIA GPUs (RTX A6000 and A100 models). As the models are deterministic, we calculate the FIM by assuming the network output is corrupted by additive Gaussian noise, as in (Berardino et al., 2017). In this case, ${\bm{I}}({\mathbf{s}})={\bm{J}}_{f}({\mathbf{s}})^{\top}{\bm{J}}_{f}({\mathbf{s}})$, where ${\bm{J}}_{f}({\mathbf{s}})$ is the Jacobian of the model ${\bm{f}}(\cdot)$ at input ${\mathbf{s}}$.
Figure 2: Principal distortions of a set of early visual models. A) Log sensitivity ratios of the two principal distortions and two random distortions for each of the four models (Berardino et al., 2017). Models in nested order (LN is the most basic, LGN is the full model). Principal distortions (filled circles) separate the log ratios, while random distortions (hollow circles) do not. B) Natural image $\mathbf{s}$ and corresponding optimized principal distortions $\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}$. C) Natural image corrupted by principal distortions, with each pair scaled so as to be equally detectable by the corresponding model (abbreviations above). Models are ordered by the log ratio of their sensitivities (panel A). If a model’s thresholds are comparable to human thresholds, the two scaled distortions should be equally visible in the top and bottom images. Note: Distorted images are best viewed at high resolution.
@@ 4.1 Early visual models
As these models were explicitly trained to predict human distortion thresholds, we provide a qualitative comparison of each model’s sensitivities to human distortion sensitivity (Fig. 2C). We adjusted the relative scaling of the principal distortion so as to be equally detectable by that model while constraining the sum of the Euclidean norms of the two distortions to be a fixed value of 100; that is, we chose positive scalars $c_{1},c_{2}$ such that $d(c_{1}{\bm{\epsilon}}_{1})=d(c_{2}{\bm{\epsilon}}_{2})$ and $c_{1}+c_{2}=100$. If a model’s thresholds are comparable to human thresholds, then these rescaled distortions should be equally detectable when added to the image ${\mathbf{s}}$. Visual inspection of these images reveals that both distortions are visible when rescaled for the LGN model and the LN model, suggesting that these models are closest to human distortion thresholds. For LG, the scaled ${\bm{\epsilon}}_{2}$ distortion is not visible, while the scaled ${\bm{\epsilon}}_{1}$ distortion is immediately apparent, suggesting a strong mismatch with human observer thresholds. The same is true of the LGG model, with the roles of the two distortions swapped. These qualitative observations are consistent with the results of (Berardino et al., 2017), in which experiments on eigen-distortions suggested that the LGN model was the best of these models in terms of consistency with human distortion sensitivity. Analogous perceptual experiments could be used to quantify the visibility of the principal distortions arising from our analysis.
@@
An advantage of our framework is that it can dramatically reduce the number of distortions that are needed to differentiate a set of models. For instance, to judge how well these models of the early visual system capture human perceptual discrimination thresholds, Berardino et al. (2017) computed the extremal eigen-distortions for each of the four models, and then measured human perceptual discrimination thresholds to all eight distortions. In general, their method requires assessing visibility of $2N$ distortions, for $N$ models. Zhou et al. (2023) considered a pair of generalized eigen-distortions for each pair of models, for a total of $N(N+1)$ distortions. In contrast, our method always selects the two distortions that maximize the variance across the models, independent of $N$. Human sensitivities to these distortions can then be estimated in perceptual discrimination experiments to judge which model(s) are closest in terms of the metric we defined Equation 3. The models whose sensitivities are far from human sensitivities can be discarded and this procedure can be repeated to best differentiate the remaining models, and so on. If one could reduce the number of models by, say, a factor of two on each iteration of this process, the total number of stimuli to be assessed scales as $2\log_{2}(N)$. This dramatic improvement in efficiency could enable the comparison of significantly larger sets of models than feasible with previous methods.
@@ 4.2 Deep neural networks
Figure 3: AlexNet versus ResNet50. A) Example base images. B) Optimized principal distortions (scaled by a factor of 20 for visibility, and using the convention $\|{\bm{\epsilon}}\|=1$ here and in other figures). C) Base image plus principal distortions. D) Log sensitivity ratios of principal distortions when comparing image representations at multiple layers of in AlexNet and ResNet50. Assignment of ${\bm{\epsilon}}_{1}$ and ${\bm{\epsilon}}_{2}$ was chosen so that the final tested layer of AlexNet has a positive log ratio. E) Log sensitivity ratios averaged across 100 base images (error bars are standard deviation). The principal distortions organize the networks by architecture—the log sensitivity ratios of AlexNet and ResNet50 are separated and early layers have smaller log ratios than late layers. AlexNet is more sensitive to distortion ${\bm{\epsilon}}_{1}$, which is concentrated on higher contrast or textured parts of the image (often the foreground object). ResNet50 is more sensitive to distortion ${\bm{\epsilon}}_{2}$ which concentrates power on relatively smooth parts of the image, such as regions of constant intensity/color.
@@
Deep Neural Networks (DNNs), originally developed for object recognition, have also been examined as models of the primate visual system (Yamins & DiCarlo, 2016; Schrimpf et al., 2018; Lindsay, 2021). A plethora of models, varying in architecture and training techniques, have been proposed, but many of these models perform quite similarly on behavioral tasks or neural benchmarks (Schrimpf et al., 2018; Tuckute et al., 2023; Conwell et al., 2022). This situation offers a well-aligned opportunity for use of our principal distortion method.
We first measured the FIM of a set of layers from two different architectures trained on the ImageNet object classification task—AlexNet (Krizhevsky et al., 2012) and ResNet50 (He et al., 2016)—and generated the principal distortions that maximally separate these models (Fig. 3, see Sec. D of the supplement for layer choices and model details). Although these architectures are not currently state-of-the-art at image recognition or neural prediction, they have been widely used and trained with various optimization strategies. Notably, the hierarchical structure of the models is reflected in the log ratios of the sensitivities, where early layers of the models are closer together in the metric space (Fig. 3D), and late layers of AlexNet are always more sensitive to ${\bm{\epsilon}}_{1}$ when late layers of ResNet50 are more sensitive to ${\bm{\epsilon}}_{2}$. There is additional structure revealed by these principal distortions—AlexNet is more sensitive to the principal distortion that generally appears concentrated on regions of the image that have more variability, i.e., the “stuff” in the image (distortion ${\bm{\epsilon}}_{1}$ in Fig. 3B), while ResNet50 is more sensitive to distortions that occur in the relatively constant regions of the image (distortion ${\bm{\epsilon}}_{2}$ in Fig. 3B). The separability of the models and the sensitivity of the networks to the two distortion types was remarkably consistent across a set of 100 base images chosen from the ImageNet dataset (Fig. 3E, see Fig. 6 of the supplement for additional examples). This distinction was also found in images designed explicitly to test for possible differences between models due to edge artifacts or contrast sensitivity (Fig. 7 of the supplement). As far as we know, this qualitative difference in sensitivities of the architectures to distortions located at different places of the image has not been documented, demonstrating that our method can reveal interpretable differences in the local sensitivities of complex computational models.
@@
The architectural difference observed between ImageNet-trained AlexNet and ResNet50 suggested that the texture of the image may be driving some of the differences in local geometry. Previous work demonstrated that standard DNNs exhibit strong “texture bias” (Geirhos et al., 2019) and proposed models that explicitly reduce the texture bias by training on Stylized ImageNet (SIN), a set of images that retains the content of each ImageNet image but overlays details matched to particular texture. Training on these images can reduce the model’s reliance on texture for classification. If this training set strongly affected the local geometry of the networks, we might expect that principal distortions generated for a mixture of architectures and training sets would be driven by this training set distinction. We find evidence that this is not the case. We generated principal distortions for 100 base images using a set of layers from two architectures (ResNet50 and AlexNet) and two different training datasets (ImageNet and Stylized ImageNet) (Fig. 4A). The principal distortions appeared qualitatively similar to those generated when we investigated only the standard networks (Fig. 4B, more examples in Fig. 8 of the supplement). Both AlexNet architectures, regardless of the training type, were more sensitive to the perturbations of higher variability parts of the image, while both ResNet50 models were more sensitive to the distortions that were mainly targeting constant areas of the image.
@@
Another well-known example of the use of training set modifications to achieve robustness in object recognition is that of adversarial training (AT). Adversarial examples are generated at each step of model training and these stimuli are used to update the model weights, with the “true” category label used for the update. Previous work has found that adversarially-trained networks are more aligned with those of biological systems (Madry, 2017; Feather et al., 2023; Gaziv et al., 2023). As adversarial examples are constrained to be very small perturbations, it seems plausible that the local geometry for AT models would differ from their standard counterparts. Indeed, we see that the principal distortions generated from the set of models that included standard and AT trained AlexNet and ResNet50 models reliably separate the model classes by training type, rather than architecture (Fig. 5A). Most layers of the adversarially trained models are more sensitive to relatively smooth changes of patches of color in the image, or to shading around edges, while most layers of the standard models are more sensitive to what appears as unstructured noise (Fig. 5B, additional examples in Fig. 9 of the supplement). This is in stark contrast to the style-trained networks of the previous section, for which principal distortions reliably separated the models by architecture. These examples provide a compelling demonstration that our method can be used to separate collections of similar models, and points to its utility in probing complex high-level representations.
@@ 5 Discussion
We have introduced a metric on image representations that quantifies differences in local geometry, and used it to synthesize “principal distortions” that maximize the variance of the metric over a set of models. When applied to hand-engineered models of the early visual system and to deep neural networks, our approach produced novel distortions for distinguishing the corresponding models. In particular, with the DNN analysis, we revealed that there are qualitative differences in local geometries of ResNet50 and AlexNet architectures, that adversarial training dramatically changes the local geometry, but stylized ImageNet training does not. Although our qualitative examples in this targeted set of neural networks do not fully elucidate the recent observations that many different models are equally good at capturing brain representations, our method provides a direct approach to begin to tease apart the interplay between local geometry and global structure.
@@
There are several limitations in our formulation of principal distortions. First, our framework is based on local differential analyses of a model at a base image, so the sensitivity estimates we obtain via these analyses can only be guaranteed to hold in an infinitesimally small neighborhood of the base image. If a model is highly nonlinear in the vicinity of a base image, then the local linear approximation may not accurately reflect model sensitivities. Second, to compute the FIM of a deterministic model, we assume additive Gaussian response noise, which is not generally representative of neural responses in the brain. Poisson variability could be used to better capture neural responses. Another approach is to fit the model noise structure to measurements of a neural system; see (Ding et al., 2023) for work along these lines.
@@
The results with DNNs reveal some intriguing properties: some models have stronger sensitivity biases in the local geometry for perturbing higher variability regions of the space, while others have more sensitivity to perturbations in relatively blank regions (see Fig. 7 of the supplement for direct evidence of this). This also provides and interesting question for future distortion detection: in what contexts are humans more sensitive to perturbations of the “stuff” of the image compared to perturbations in empty parts of the image? And how do these observed differences relate to previous work investigating how human and neural networks rely on on spatial frequencies for classification (Subramanian et al., 2023)? Finally, beyond direct comparisons with human observers using psychophysics experiments, our method may be useful in the domain of neural network interpretability, where it may be useful to have direct comparisons of the local distortions that will maximally differentiate sets of models.
Acknowledgements

### v2

@@ 3.2 Principal distortions for comparing multiple image representations
$\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}=\operatorname*{arg\,max}_{{\mathbf{u}},{\mathbf{v}}}\sum_{n=1}^{N}\sum_{m=1}^{N}m_{{\mathbf{u}},{\mathbf{v}}}^{2}({\bm{I}}_{A_{n}},{\bm{I}}_{A_{m}}).$ (4)
@@
$\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}=\operatorname*{arg\,max}_{{\mathbf{u}},{\mathbf{v}}}\sum_{n=1}^{N}\left|\log\frac{d_{A_{n}}({\mathbf{u}})}{d_{A_{n}}({\mathbf{v}})}-\frac{1}{N}\sum_{m=1}^{N}\log\frac{d_{A_{m}}({\mathbf{u}})}{d_{A_{m}}({\mathbf{v}})}\right|^{2}$
@@
There are several other natural extensions of generalized-eigendistortions when considering $N>2$ models. For example, for any $p\geq 1$, one can choose distortions ${\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}$ that maximize the sum of the $p$th power of all pairwise distances. Here we focus on the case of maximizing the variance ($p=2$) and leave an exploration of other moments to future work.
@@ 4 Experimental results
As a demonstration of our method, we generated principal distortions for computational models previously proposed to capture aspects of the human visual system. All models were implemented in PyTorch (Ansel et al., 2024) and simulations were performed on NVIDIA GPUs (RTX A6000 and A100 models). As the models are deterministic, we calculate the FIM by assuming the network output is corrupted by additive Gaussian noise, as in (Berardino et al., 2017). In this case, ${\bm{I}}({\mathbf{s}})={\bm{J}}_{f}({\mathbf{s}})^{\top}{\bm{J}}_{f}({\mathbf{s}})$, where ${\bm{J}}_{f}({\mathbf{s}})$ is the Jacobian of the model ${\bm{f}}(\cdot)$ at input ${\mathbf{s}}$ and the induced geometry on stimulus space is the pullback via ${\bm{f}}(\cdot)$ of the Euclidean geometry on representation space.
Figure 2: Principal distortions of four early visual models. A) Log sensitivity ratios of the two principal distortions and two random distortions for each model (Berardino et al., 2017). Models are nested (LN is the most basic, LGN is the full model). Principal distortions (filled circles) give rise to diverse log ratios, while random distortions (hollow circles) do not. B) Natural image $\mathbf{s}$ and corresponding optimized principal distortions $\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}$. C) Natural image corrupted by principal distortions, with each pair scaled so as to be equally detectable by one model (as indicated above). Models are ordered by the log ratio of their sensitivities (panel A). If a model’s thresholds are proportional to human thresholds, the corresponding pair of scaled distortions should be equally visible in the top and bottom images. Note: Images are best viewed at high resolution.
@@ 4.1 Early visual models
As these models were explicitly trained to predict human distortion thresholds, we provide a qualitative comparison of each model’s sensitivities to human distortion sensitivity (Fig. 2C). For each model, we adjusted the relative scaling of each principal distortion until the model was equally sensitive to the scaled distortions while constraining the sum of the Euclidean norms of the two distortions to be a fixed value of 100; that is, we chose positive scalars $k_{1},k_{2}$ such that the sensitivities were equal, $d(k_{1}{\bm{\epsilon}}_{1})=d(k_{2}{\bm{\epsilon}}_{2})$, with $k_{1}+k_{2}=100$. If a model’s thresholds are comparable to human thresholds, then the pair of images should also appear equally distorted to a human observer. Visual inspection of these images reveals that both distortions are visible when rescaled for the LGN model and the LN model, suggesting that these models are closest to human distortion thresholds. For LG, the scaled ${\bm{\epsilon}}_{2}$ distortion is not visible, while the scaled ${\bm{\epsilon}}_{1}$ distortion is immediately apparent, suggesting a strong mismatch with human observer thresholds. The same is true of the LGG model, with the roles of the two distortions swapped. These qualitative observations are consistent with the results of (Berardino et al., 2017), in which experiments on eigen-distortions suggested that the LGN model was the best of these models in terms of consistency with human distortion sensitivity. Future work with analogous human perceptual experiments could directly quantify the visibility of the principal distortions arising from our analysis (see Supp. Fig. SI.1 for an illustration).
@@
An advantage of our framework is that it can dramatically reduce the number of distortions needed to differentiate a set of models. For instance, to judge how well these models of the early visual system capture human perceptual discrimination thresholds, Berardino et al. (2017) computed the extremal eigen-distortions for each of the four models, and then measured human perceptual discrimination thresholds to all eight distortions. In general, their method requires assessing visibility of $2N$ distortions, for $N$ models. Zhou et al. (2023) considered a pair of generalized eigen-distortions for each pair of models, for a total of $N(N+1)$ distortions. In contrast, our method always selects the two distortions that maximize the variance across the models, independent of $N$. Human sensitivities to these distortions can then be estimated in perceptual discrimination experiments to judge which model(s) are closest in terms of the metric we defined Equation 3. The models whose sensitivities are far from human sensitivities can be discarded and this procedure can be repeated to best differentiate the remaining models, and so on. If one could reduce the number of models by, say, a factor of two on each iteration of this process, the total number of stimuli to be assessed scales as $2\log_{2}(N)$. This dramatic improvement in efficiency could enable the comparison of significantly larger sets of models than feasible with previous methods.
@@ 4.2 Deep neural networks
Figure 3: AlexNet versus ResNet50. A) Example base images. B) Optimized principal distortions (scaled by a factor of 20 for visibility, and using the convention $\|{\bm{\epsilon}}\|=1$ here and in other figures). C) Base image plus principal distortions. D) Log sensitivity ratios of principal distortions when comparing image representations at multiple layers of AlexNet and ResNet50. Assignment of ${\bm{\epsilon}}_{1}$ and ${\bm{\epsilon}}_{2}$ was chosen so that the final tested layer of AlexNet has a positive log ratio. E) Log sensitivity ratios averaged across 100 base images (error bars are standard deviation). The principal distortions organize the networks by architecture—the log sensitivity ratios of AlexNet and ResNet50 are separated and early layers have smaller log ratios than late layers. AlexNet is more sensitive to distortion ${\bm{\epsilon}}_{1}$, which is concentrated on higher contrast or textured parts of the image (often the foreground object). ResNet50 is more sensitive to distortion ${\bm{\epsilon}}_{2}$ which concentrates power on relatively smooth parts of the image, such as regions of constant intensity/color.
@@
Deep Neural Networks (DNNs), originally developed for object recognition, have also been examined as models of the primate visual system (Yamins & DiCarlo, 2016; Schrimpf et al., 2018; Lindsay, 2021). A plethora of models, varying in architecture and training techniques, have been proposed, but many of these models perform quite similarly on behavioral tasks or neural benchmarks (Schrimpf et al., 2018; Tuckute et al., 2023; Conwell et al., 2024). This situation offers a well-aligned opportunity for use of our principal distortion method. Here, we investigate previously trained sets of models (no additional training is required to apply our method) and generate principal distortions to differentiate the image representations.
We first investigated a set of layers from two architectures trained on the ImageNet object classification task—AlexNet (Krizhevsky et al., 2012) and ResNet50 (He et al., 2016)—and generated the principal distortions that maximally separate these models (Fig. 3, see Sec. D of the supplement for layer choices and model details). Although these architectures are not currently state-of-the-art at image recognition or neural prediction, they have been widely used and trained with various optimization strategies, making them well suited for controlled experiments. (See Supp. Fig. SI.2 for a similar test of the method on the modern architectures EfficientNet and ViT.)
Notably, the hierarchical structure of the models is reflected in the log ratios of the sensitivities, where early layers of the models are closer together in the metric space (Fig. 3D), and late layers of AlexNet are always more sensitive to ${\bm{\epsilon}}_{1}$ when late layers of ResNet50 are more sensitive to ${\bm{\epsilon}}_{2}$. There is additional structure revealed by these principal distortions—AlexNet is more sensitive to the principal distortion that generally appears concentrated on regions of the image that have more variability, i.e., the “stuff” in the image (distortion ${\bm{\epsilon}}_{1}$ in Fig. 3B), while ResNet50 is more sensitive to distortions that occur in the relatively constant regions of the image (distortion ${\bm{\epsilon}}_{2}$ in Fig. 3B). The separability of the models and the sensitivity of the networks to the two distortion types was remarkably consistent across a set of 100 base images chosen from the ImageNet dataset (Fig. 3E, see Supp. Fig. SI.3 for additional examples). These distortions differentially affected the classification decisions of the two architectures (Supp. Fig. SI.4). The distortions also replicated in a set of six models (3 randomly initialized AlexNet models and 3 randomly initialized ResNet50 models, each trained on ImageNet-1k, see Supp. Fig. SI.5). This distinction was also found in images designed explicitly to test for possible differences between models due to edge artifacts or contrast sensitivity (Supp. Fig. SI.6). As far as we know, this qualitative difference in sensitivities of the architectures to distortions in different portions of the image has not been documented, demonstrating that our method can reveal interpretable differences in the local sensitivities of complex computational models.
@@
The architectural difference observed between ImageNet-trained AlexNet and ResNet50 suggested that the texture of the image may be driving some of the differences in local geometry. Previous work demonstrated that standard DNNs exhibit strong “texture bias” (Geirhos et al., 2019) and proposed models that explicitly reduce the texture bias by training on Stylized ImageNet (SIN), a set of images that retains the content of each ImageNet image but overlays details matched to particular texture. Training on these images can reduce the model’s reliance on texture for classification. If this training set strongly affected the local geometry of the networks, we might expect that principal distortions generated for a mixture of architectures and training sets would be driven by this training set distinction. We find evidence that this is not the case. We generated principal distortions for 100 base images using a set of layers from two architectures (ResNet50 and AlexNet) and two different training datasets (ImageNet and SIN, Fig. 4A). The principal distortions appeared qualitatively similar to those generated when we investigated only the standard networks (Fig. 4B, more examples in Supp. Fig. SI.7). Both AlexNet architectures, regardless of the training type, were more sensitive to the perturbations of higher variability parts of the image, while both ResNet50 models were more sensitive to the distortions that were mainly targeting constant areas of the image. To quantify this observation and dissociate the role of foreground and background information from the high- and low-frequency content in the image, we ran an experiment using the ImageNet-9 dataset with foreground and background masks, where we generated principal distortions after applying a low-pass filter to either the foreground or the background of the image (Supp. Fig. SI.8). The principal distortion that AlexNet is more sensitive to consistently concentrated on portions of the image with higher spatial frequencies (i.e., parts of the image that are not smoothed), while ResNet50 is concentrated on parts of the image with only low-frequency information (the blurred parts of the image, regardless of whether this is in the foreground or background).
@@
Another well-known example of the use of training set modifications to achieve robustness in object recognition is that of adversarial training (AT). Adversarial examples are generated at each step of model training and these stimuli are used to update the model weights, with the “true” category label used for the update. Previous work has found that adversarially-trained networks are more aligned with those of biological systems (Madry, 2017; Feather et al., 2023; Gaziv et al., 2023). As adversarial examples are constrained to be very small perturbations, it seems plausible that the local geometry for AT models would differ from their standard counterparts. Indeed, we see that the principal distortions generated from the set of models that included standard and AT trained AlexNet and ResNet50 models reliably separate the model classes by training type, rather than architecture (Fig. 5A). Most layers of the adversarially trained models are more sensitive to relatively smooth changes of patches of color in the image, or to shading around edges, while most layers of the standard models are more sensitive to what appears as unstructured noise (Fig. 5B, additional examples in Supp. Fig. SI.9). This is in stark contrast to the SIN-trained networks of the previous section, for which principal distortions reliably separated the models by architecture. These examples demonstrate that our method can be used to separate collections of similar models, and points to its utility in probing complex high-level representations.
@@ 5 Discussion
We have introduced a metric on image representations that quantifies differences in local geometry, and used it to synthesize “principal distortions” that maximize the variance of the metric over a set of models. When applied to hand-engineered models of the early visual system and to DNNs, our approach produced novel distortions for distinguishing the corresponding models. In particular, with the DNN analysis, we revealed that there are qualitative differences in local geometries of ResNet50 and AlexNet architectures, that adversarial training dramatically changes the local geometry, but stylized ImageNet training does not. Although our qualitative examples in this targeted set of neural networks do not fully elucidate the recent observations that many different models are equally good at capturing brain representations, our method provides a direct approach to begin to tease apart the interplay between local geometry and global structure.
@@
There are several limitations in our formulation of principal distortions. First, our framework is based on local differential analyses of a model at a base image, so the sensitivity estimates we obtain via these analyses can only be guaranteed to hold in an infinitesimally small neighborhood of the base image. If a model is highly nonlinear in the vicinity of a base image, then the local linear approximation may not accurately reflect model sensitivities. Second, to compute the FIM of a deterministic model, we assume additive Gaussian response noise, which is not generally representative of neural responses in the brain. Poisson variability could be used to better capture neural responses. Another approach is to fit the model noise structure to measurements of a neural system; see (Ding et al., 2023; Nejatbakhsh et al., 2023) for work along these lines.
@@
The results with DNNs reveal some intriguing properties: some models have stronger sensitivity biases in the local geometry for perturbing higher frequency regions of the space, while others have more sensitivity to perturbations in relatively low frequency regions. This also suggests an interesting question for future distortion detection: in what contexts are humans more sensitive to perturbations of the “stuff” of the image compared to perturbations in empty parts of the image? And how do these observed differences relate to previous work investigating how human and neural networks rely on spatial frequencies for classification (Subramanian et al., 2023)? Finally, beyond direct comparisons with human observers using psychophysics experiments, our method may be useful in the domain of neural network interpretability, where it may be useful to have direct comparisons of the local distortions that will maximally differentiate sets of models. Principal distortions are complementary to many currently used model interpretability tools (e.g., saliency maps, attention weight visualization) since they are explicitly designed to generate distortions that differentiate a set of models, as opposed to interpreting one model at a time.
Acknowledgments